OpenStudy (anonymous):

What is the 4th root of 8 multiplied by the 4th root of 10?

OpenStudy (anonymous):

\[\sqrt[n]a\times\sqrt[n]b=\sqrt[n]{ab}\]

OpenStudy (anonymous):



OpenStudy (anonymous):